Grade 12th passAlgebra

Find the ratio of
₹ 5 to 50 paise
Find the ratio of
₹ 5 to 50 paise

Harsh Singh Chauhan
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We know that,
₹1 = 100 paise
Then,
₹5 = 5 × 100 = 500 paise
Now we have to find the ratio,
= 500/50
= 10/1
So, the required ratio is 10: 1

Thanks